### v3

# MVP-Shapley: Feature-based Modeling for Evaluating the Most Valuable Player in Basketball

Abstract
The burgeoning growth of the esports and multiplayer online gaming community has highlighted the critical importance of evaluating the Most Valuable Player (MVP). The establishment of an explainable and practical MVP evaluation method is very challenging. In our study, we specifically focus on play-by-play data, which records related events during the game, such as assists and points. We aim to address the challenges by introducing a new MVP evaluation framework, denoted as MVP-Shapley, which leverages Shapley values. This approach encompasses feature processing, win-loss model training, Shapley value allocation, and MVP ranking determination based on players’ contributions. Additionally, we optimize our algorithm to align with expert voting results from the perspective of causality. Finally, we substantiated the efficacy of our method through validation using the NBA dataset and the Dunk City Dynasty dataset and implemented online deployment in the industry. Our code is available at https://anonymous.4open.science/r/MVP-Shapley-F18B/ and our data is available at https://anonymous.4open.science/r/nba-data-1078.
I Introduction
As the esports and multiplayer online gaming community continues to expand, recognizing the Most Valuable Player (MVP) holds substantial importance. The MVP is often the individual who has made the most significant contributions to their team during competitions, showcasing exceptional skills, tactics, and teamwork abilities. Beyond enhancing an individual’s prestige, MVP evaluation serves as a catalyst, inspiring players to become core team members, thus elevating overall gameplay and teamwork standards. This form of acknowledgment encourages players to pursue gaming excellence and fosters a more competitive and interactive gaming environment. Therefore, developing a fair MVP evaluation method is paramount for the healthy progression of the gaming community and the esports industry.
Figure 1: There are various player statistics in play-by-play data; the challenge is determining how to utilize these data to evaluate the MVP.
In the context of basketball games, data can be categorized into two distinct types: (i) tracking data, which is usually collected using optical or device tracking and processing systems to capture the movement and trajectory of players or the ball on the court [1]; (ii) play-by-play data, which records a series of related events that occur during the game, such as assists and points. However, tracking data is generally more expensive and difficult to obtain [2], so we mainly analyze play-by-play data. Fig. 1 illustrates that the player statistics described in TABLE II are multidimensional and complex. Determining how to effectively use these data to evaluate MVPs is a topic worth exploring.
TABLE I: Comparison with Existing Works.
Method Scalability Explainability Collinearity Issue High Overhead Costs
Metric Weighting $\times$
Voting Selection $\times$
Machine Learning Techniques $\times$
Cooperative Game Theory among Players $\times$ $\times$ $\times$
Ours $\times$ $\times$
  • 1. , $\checkmark$: Yes; , $\times$: No; : Partial, which means a significant gap exists between the current state and the desired goal.
  • 2. : Applicable to evaluating single-game MVP, regular season MVP, and other games with play-by-play data.
  • 3. : The cost includes manpower, material resources, and money.
The existing methods for evaluating MVP are summarized into four categories: Metric Weighting, Voting Selection, Machine Learning Techniques, and Cooperative Game Theory among Players. Metric weighting methods encompass a variety of single-metric evaluations, such as PM, APM [3, 4], BPM [5, 6], RPM [7], WS, WS48 [8], WARP, VORP [9], among others. The weighting of different metrics was introduced early on [10, 11]. There are also numerous industry applications. Typically, game strategists, drawing from experience, devise weighted formulas incorporating various metrics to determine the MVP. For example, in Netease’s "Dunk City Dynasty", different metrics such as points, assists, rebounds, steals, and blocks are assigned weights to compute the MVP scores. Tencent’s "Honor of Kings" employs a formula based on KDA data. The metric weighting methods lack interpretability, as weights are arbitrarily set based on experience, potentially leading to high-weighted metrics not necessarily contributing significantly to the win rate. Additionally, there may be collinearity among metrics, making MVP calculation challenging. For voting selection, in competitions such as the NBA, experienced experts are tasked with voting for the MVP based on players’ game statistics, relying on their observations and judgment. This paper assumes that voting is fair, just, and transparent, and uses the final voting results as the ground truth. However, this method is costly due to the significant human resources required for voting. The rise of artificial intelligence has led to the adoption of various machine-learning techniques in MVP evaluation. Many of these methods [12, 13, 14, 15, 16] use black box models to predict player value and ignore the collinearity issues between players. The cooperative game theory approach used to analyze the origins of sports [17, 18] can also be employed to assess the value of basketball players. Metulini et al. [19] were the first to apply the Shapley value to player value calculation, identifying the most valuable player by considering different team lineup combinations. However, their approach is limited to assessing the value of players within a single, stable team, and is difficult to extend to the entire league. Furthermore, this method is not applicable for evaluating the MVP in an online basketball game, where the team formation may be random and unstable.
Our study faces two main challenges: 1). Online games with random team formations have limited historical combinations of players, making it difficult to apply the leave-one-out technique for players by traditional Shapley methods. 2). Player statistics are complex and diverse, with the presence of confounding variables, which complicates the evaluation process.
To address these challenges, we propose an MVP evaluation method based on Shapley values, referred to as MVP-Shapley, which consists of five key steps: feature processing and evaluation, training the win-loss model, Shapley value allocation, MVP evaluation, and algorithmic refinements. Specifically, MVP-Shapley uses a win-loss model, trained on players’ statistical feature data, as the utility function in the Shapley value formula. The Shapley value formula then calculates each player’s contribution to the team’s wins or losses. Based on these calculated contributions, MVP rankings are determined. Finally, the algorithm is optimized by incorporating causal inference techniques to reduce the impact of confounding variables. Compared to existing work, our advantages are summarized in TABLE I. The main contributions are summarized as follows:
$\bullet$ We introduce MVP-Shapley, a Shapley value-based framework for evaluating the MVP, which is highly interpretable and scalable. This approach differs from traditional player-centric evaluation by emphasizing the impact of individual features.
$\bullet$ To enhance the algorithm and better align it with true MVP voting, we analyze the optimal combination of variables from a causal perspective. We introduce an optimization process to reduce the complexity of variable combinations, along with a technique to fuzzify confounding variables.
$\bullet$ For the NBA dataset, we have compiled an extensive dataset of all regular-season and finals NBA games since 2000, incorporating a comprehensive array of both basic and advanced player statistics. Additionally, we leverage crowdsourcing to gather MVP rankings for the Dunk City Dynasty dataset through voting, using these rankings as the ground truth to validate our method.
$\bullet$ We conducted thorough experiments on the NBA and Dunk City Dynasty datasets and implemented online deployment in the industry, comparing our method to existing approaches. The results demonstrate that our method outperforms the others.
II Preliminaries
TABLE II: NBA Statistics Summary
Basic Statistics Description
MP Minutes Played
FG Field Goals Made
FGA Field Goals Attempted
FG% Field Goal Percentage
3P 3-Point Field Goals Made
3PA 3-Point Field Goals Attempted
3P% 3-Point Field Goal Percentage
FT Free Throws Made
FTA Free Throws Attempted
FT% Free Throw Percentage
ORB Offensive Rebounds
DRB Defensive Rebounds
TRB Total Rebounds
AST Assists
STL Steals
BLK Blocks
TOV Turnovers
PF Personal Fouls
PTS Points
+/- Plus/Minus
Advanced Statistics Description
TS% True Shooting Percentage
eFG% Effective Field Goal Percentage
3PAr 3-Point Attempt Rate
FTr Free Throw Attempt Rate
ORB% Offensive Rebound Percentage
DRB% Defensive Rebound Percentage
TRB% Total Rebound Percentage
AST% Assist Percentage
STL% Steal Percentage
BLK% Block Percentage
TOV% Turnover Percentage
USG% Usage Percentage
ORtg Offensive Rating
DRtg Defensive Rating
BPM Box Plus/Minus
In this work, we focus on the MVP evaluation problem using play-by-play data and Shapley values [20]. In this section, we will introduce the datasets and Shapley values.
II-A Datasets
II-A1 NBA Dataset
We crawled all regular season and playoff play-by-play data of the NBA since 2000 from the basketball reference website ( www.basketball-reference.com/). There are 31,627 games in total, and four tables for each of the two teams in each game: one basic statistics table and one advanced statistics table for each team. Each game contains 20 basic statistics and 15 advanced statistics of all players on the court. TABLE II shows the description of basic and advanced statistics. Our collated dataset is available at https://anonymous.4open.science/r/nba-data-1078.
II-A2 Dunk City Dynasty Dataset
Dunk City Dynasty  [^1] https://www.dunkcitymobile.com/ is a 3V3 basketball competitive mobile game developed by Netease and authorized by the NBA official players union. The dataset comprises play-by-play data, including the following 82 statistics such as ‘BeStrongDisturbedShoot’, ‘NicePass’, ‘BuffUltraSkillSteal’, ‘TwoShots’, ‘Block’, ‘DisturbedLayupNum’, and others. A total of 184,908 games are selected for the Summit match on March 15, 2024.
II-B Shapley Values
The Shapley value (SV) [20] is commonly used in game theory to identify the contributions of players collaborating in a coalition. Assume there are $n$ players $N=\{1,\cdots,n\}$, $S\subset N$ is a subset of this coalition. Given a utility function $v:S\to\mathbb{R}$, the SV of the player $i$ is defined as $\phi_{i}(v)$, which is the average marginal contribution of $i$ to all possible subsets of $S$:
$\phi_{i}(v)=\frac{1}{n}\sum_{S\subset N/\{i\}}{\tbinom{n-1}{|S|}}^{-1}(v(S\cup\{i\})-v(S)).$ (1)
Here, $|S|$ represents the size of the subset $S$. The term $\frac{1}{n}{\tbinom{n-1}{|S|}}^{-1}$ represents the probability of the subset $S$ appearing. The term $v(S\cup i)-v(S)$ represents the marginal contribution of player $i$ to the sub-coalition $S$.
III Methods
Our core idea involves leveraging play-by-play data statistics to develop a model that predicts win-loss outcomes. Subsequently, we utilize this model to assess individual player contributions via feature-level Shapley value analysis, ultimately evaluating MVPs. The methodology is structured into five distinct phases: (1) feature preprocessing, (2) win-loss prediction model development, (3) computation of players’ contributions, and (4) MVP evaluation. (5) Finally, to improve the algorithm’s alignment with empirical voting results, we refine our approach from a causal perspective by incorporating feature combinations and fuzzification.
III-A Feature Preprocessing
[…]
$x_{1}$ $=(a_{11},\cdots,a_{1q},\cdots,a_{p1},\cdots,a_{pq},\cdots,b_{11},\cdots,$ (2)
$b_{1q},\cdots,b_{p1},\cdots,b_{pq}),\ y_{1}=1;$
$x_{2}$ $=(b_{11},\cdots,b_{1q},\cdots,b_{p1},\cdots,b_{pq},\cdots,a_{11},\cdots,$
$a_{1q},\cdots,a_{p1},\cdots,a_{pq}),\ y_{2}=0.$
The dataset for $m$ games is $\{(x_{i},y_{i})|i\in\{1,\cdots,2m\}\}$.
III-B Win-Loss Prediction Model Development
After constructing the dataset, we select $n$ data points as the training set. We employ the LightGBM model [21] for predictive analysis. LightGBM is a gradient-boosting framework that utilizes tree-based learning algorithms and is engineered for efficiency and high performance. The model minimizes the following objective function:
$\mathcal{L}(\theta)=\sum_{i}^{n}l(y_{i},f(x_{i}))+\sum_{i}\Omega(f_{i})+\sum_{k}\Omega(\gamma_{k}),$ (3)
where $\mathcal{L}(\theta)$ denotes the overall objective function and $\theta$ denotes the model parameters. $y_{i}$ represents the true label, $f(x_{i})$ is the predicted score. $f(x_{i})=(f(x_{i})_{loss},f(x_{i})_{win})\in R^{2}$, where $f(x_{i})_{loss}+f(x_{i})_{win}=1$, and $f(x_{i})_{win}$ represents the probability of winning. $l(y_{i},f(x_{i}))$ denotes the binary log loss function and is defined as:
$l(y_{i},f(x_{i}))$ $=-(y_{i}\log(f(x_{i})_{win})+(1-y_{i})$ (4)
$\log(1-f(x_{i})_{win})).$
$\Omega(f_{i})$ denotes the regularization term for the $i$-th tree, and $\Omega(\gamma_{k})$ represents the regularization term for the $k$-th leaf. The objective function $\mathcal{L}(\theta)$ encapsulates the model’s optimization goal, aiming to minimize this function through the learning process.
III-C Computation of Players’ Contributions
MVP-Shapley first computes the Shapley values for the player’s features, then uses these values to assess the player’s contribution to the match outcome. Inspired by probability-based Shapley Value [22], the winning probability from the win-loss model $f$ is used as the utility function $v$ in the Shapley formula (1):
$v(S)=f(x_{S})_{win}.$ (5)
For a game data $x=(x^{1},\cdots,x^{2pq})\in R^{2pq}$, $S$ represents a subset of feature indices, and $x_{S}$ denotes the feature subset of $x$ corresponds to the indices in $S$. MVP-Shapley calculates the average marginal contribution $\phi_{i}$ of each feature $x^{i}$ to the winning score $f(x)_{win}$ across all possible combinations of other features, i.e., the Shapley value, and uses this as the contribution of feature $x^{i}$ to the winning rate. Let $N=\{1,\cdots,2pq\}$. The calculation of $\phi_{i}(x)$ is as follows:
$\phi_{i}(x)=\frac{1}{2pq}\sum_{S\subset N/\{i\}}{\tbinom{2pq-1}{|S|}}^{-1}(f(x_{S\cup\{i\}})_{win}-f(x_{S})_{win}).$ (6)
Next, we will explain how to evaluate a player’s contribution in a single game. We construct two data $x_{1}$ and $x_{2}$ as follows:
$x_{1}$ $=(a_{11},\cdots,a_{1q},\cdots,a_{p1},\cdots,a_{pq},\cdots,b_{11},$ (7)
$\cdots,b_{1q},\cdots,b_{p1},\cdots,b_{pq})$
$x_{2}$ $=(b_{11},\cdots,b_{1q},\cdots,b_{p1},\cdots,b_{pq},\cdots,a_{11},$
$\cdots,a_{1q},\cdots,a_{p1},\cdots,a_{pq}).$
The contribution of the player $a_{i}$ or $b_{i}$ to the game outcome can be defined as the sum of all Shapley values of the player’s features on the home team minus the sum of all Shapley values of the player’s features on the away team. This can be denoted as $\Phi(a_{i},\{x_{1},x_{2}\})$ and $\Phi(b_{i},\{x_{1},x_{2}\})$, respectively.
$\Phi(a_{i},\{x_{1},x_{2}\})$ $=\sum_{j\in\mathcal{H}(a_{i})}\phi_{j}(x_{1})-\sum_{j\in\mathcal{A}(a_{i})}\phi_{j}(x_{2}),$ (8)
$\Phi(b_{i},\{x_{1},x_{2}\})$ $=\sum_{j\in\mathcal{H}(b_{i})}\phi_{j}(x_{2})-\sum_{j\in\mathcal{A}(b_{i})}\phi_{j}(x_{1}).$
Here, $\mathcal{H}(a_{i})$ represents the index set of all features of player $a_{i}$ when he is on the home team. $\mathcal{A}(a_{i})$ represents the index set of all features of player $a_{i}$ when he is on the away team. Similarly, $\mathcal{H}(b_{i})$, $\mathcal{A}(b_{i})$ denote the index sets of all features of player $b_{i}$ when they are on the home and away teams, respectively.
III-D MVP Evaluation
III-D1 MVP evaluation of a single game
For a single game, generally speaking, the MVP is awarded to the player from the winning team. Since team A defeated team B, the player with the highest contribution value calculated from team A is the MVP of this game. The calculation is as follows:
$MVP=\arg\max_{a_{i}}\Phi(a_{i},\{x_{1},x_{2}\}).$ (9)
III-D2 MVP evaluation of multiple games
However, the evaluation of the regular season MVP and the finals MVP is based on the comprehensive performance of players across multiple games. To address this, MVP-Shapley provides three distinct methods for calculating the MVP across multiple games. Let $p_{i}$ represent a player who has participated in $T$ games, with $\Phi_{i}(p_{i})$ denoting the player’s contribution to the $i$-th game. As the previous section discussed, we have established a method for calculating the contribution of a player in a single game.
1) Method 1 ($M_{1}$): The MVP is determined by the player with the smallest average ranking of their contribution across all winning games they participated in. Let $G_{win}$ represent the set of winning games for player $p_{i}$. The MVP is calculated as follows:
$MVP=\arg\min_{a_{i}}\frac{1}{|G_{win}|}\sum_{i\in G_{win}}rank(\Phi_{i}(p_{i})).$ (10)
Here, $rank(\Phi_{i}(p_{i}))$ represents the ranking of player $p_{i}$’s contribution $\Phi_{i}(p_{i})$ in this game from large to small.
2) Method 2 ($M_{2}$): The MVP is determined by the player with the smallest average ranking of their contribution across all games they participated in, regardless of the outcome. This method evaluates the player’s performance uniformly across all games. The MVP is calculated as follows:
$MVP=\arg\min_{a_{i}}\frac{1}{T}\sum_{i\in\{1,\cdots,T\}}rank(\Phi_{i}(p_{i})).$ (11)
3) Method 3 ($M_{3}$): The MVP is determined by the player with the largest average contribution across all games they participated in. This method focuses on the sum of the Shapley values themselves, rather than their rankings. The MVP is calculated as follows:
$MVP=\arg\max_{a_{i}}\frac{1}{T}\sum_{i\in\{1,\cdots,T\}}\Phi_{i}(p_{i}).$ (12)
III-E Refinements in Algorithmic Approach
In our analysis, we possess the ground truth of the MVP voting data with various statistical metrics related to players. We process the statistical data into features to train a win-loss model and compute the utility function. However, we are uncertain about which features are the true causal factors leading to the ground truth results. Therefore, our approach involves seeking superior causal factors from the outcomes. Let $Y$ represent the ground truth, and $X_{1},X_{2},\ldots,X_{n}$ denote the variables related to computing the utility function before feature preprocessing. Our objective is to determine the optimal set of variables $X^{*}$ such that: $X^{*}=\underset{X}{\arg\max}\,\mathbb{E}[Y|X]$, where $\mathbb{E}[Y|X]$ denotes the conditional expectation of the ground truth given the set $X$. However, the computational complexity is excessively high with $2^{n}$ possible combinations. To address this, we propose using feature importance ranking to group features with low importance together, thereby reducing the complexity of variable combinations. For example, Fig. 6 illustrates the influence of different features on win-loss model predictions of the NBA dataset. The top two most influential variables, ‘+,-’ and ‘DRtg’ can be considered as individual variables, and all the remaining variables can be considered as a whole. Then we have only 8 possible combinations.
Figure 2: Cause and effect graph.
As shown in Fig. 2, the statistical variables capture most of the behavioral performance during the game, which contributes to the outcome. However, some variables, known as confounding variables, not only impact the result but also influence other variables. Despite this, they still exert a significant effect on the results. To mitigate their confounding impact and allow other variables to have a more accurate influence, we apply fuzzification. Specifically, we discretize the variable values using bucketing. Given a variable $X_{i}$ representing ‘DRtg’, we discretize the variable by binning the values into intervals. Let $t$ be the number of bins, and $B=\{bin_{1},bin_{2},\cdots,bin_{t}\}$ be the set of bin boundaries. The discretized variable is defined as follows:
[…]
Here, $\hat{X_{i}}$ represents the discretized variable value of ‘DRtg’ and is utilized to mitigate its impact on the determination of the MVPs. The boundaries $B$ are determined based on the distribution of ‘DRtg’ values, ensuring that the discretization process effectively reduces the influence of ‘DRtg’ on the MVP evaluation process. When $|B|=1$, it is equivalent to completely removing the variable. Fig. 6 shows the influence of the features after fuzzification ($|B|=3$), showing a more balanced distribution of their impacts.
IV Experiments
IV-A Experimental Setup
[…]
IV-A1 Baselines
1) Ground Truth (GT): NBA Dataset: The NBA regular season MVP rankings and Finals MVP awards are used as the ground truth, with detailed information available on the website:  https://www.basketball-reference.com/awards. Dunk City Dynasty Dataset: For the Dunk City Dynasty dataset, crowdsourcing is used for MVP ranking voting on the NetEase Youling Crowdsourcing Platformhttps://zb.163.com/mark/task). Annotators, who are professional basketball enthusiasts, carefully review game videos and rank the winning players (A total of $3!=6$ rankings), considering factors beyond just points, assists, defense, rebounds, steals, and turnovers, such as overall contributions to the game. Annotators should objectively choose the best one from 6 ranking options. A total of 500 game videos are included in our crowdsourcing effort. Finally, the platform employs truth inference algorithms [24, 25] to establish the confidence level of the MVP ranking for each game, termed crowdsourced confidence.
2) GSv [19]: According to the team’s historical lineup data, GSv fits different lineup win rate prediction models and uses the generalized Shapley value to calculate each team’s MVP and best lineup.
3) API [9]: This is a metric weighting method, each advanced efficiency value metric is normalized and weighted to obtain the API.
IV-A2 Evaluation Metrics
1) Average Rank Difference (ARD): The average rank difference quantifies the average discrepancy between the predicted MVP rank and the actual rank of the corresponding player. Mathematically, ARD is calculated as follows: $ARD=\frac{1}{N}\sum_{i=1}^{N}|rank_{\text{predicted MVP}_{i}}-rank_{\text{ground truth}_{i}}|.$ NN is the total number of players on the MVP voting list. $rank_{\text{predicted MVP}_{i}}$ is the rank of the predicted MVP for the i-th player on the MVP voting list, $rank_{\text{ground truth}_{i}}$ is the actual voting rank of the i-th player.
2) Spearman’s Rank Correlation Coefficient [26] (SRCC): Spearman’s Rank Correlation Coefficient is a measure of the strength and direction of association between two ranked variables. The formula for Spearman’s Rank Correlation Coefficient is given by: $SRCC=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}.$ Here, $d_{i}$ represents the difference between the ranks of corresponding players, and $n$ is the number of players. Spearman’s Rank Correlation Coefficient ranges from -1 to 1, where 1 indicates a perfect positive monotonic relationship, and -1 indicates a perfect negative monotonic relationship.
3) Recall (R): The recall here is defined as the proportion of players ranked in the top K of the MVP voting appearing in the predicted top K by one MVP evaluation method. Let $G_{K}$ represent the set of top K players in MVP voting, and $M_{K}$ represent the set of top K players predicted by the method. This can be expressed as: $R=\frac{|G_{K}\bigcap M_{K}|}{K}.$
4) Accuracy (ACC): Assuming there are $n$ matches, each with an MVP voting result. Let $m$ denote the number of matches for which one method’s evaluation for MVP matches the voting result. The accuracy can be expressed as: $\text{ACC}=\frac{m}{n}.$
IV-B Key Differences Between NBA and Dunk City Dynasty
IV-B1 Ground Truth Availability
  • NBA: Uses expert-voted MVP rankings (real-world ground truth from  https://www.basketball-reference.com/awards)
  • Dunk City Dynasty (DCD): As an online game, it lacks intrinsic ground truth. We:
    • Crowdsourced MVP rankings via NetEase’s platform
    • Employed truth inference algorithms ( [24, 25]) and the state-of-the-art voting algorithm, the GOSDT tree [27] to ensure reliability
IV-B2 Game Structure & MVP Evaluation Scope
TABLE III: Structural Comparison Between NBA and DCD
Feature NBA DCD
Evaluation Scope Multi-game performance Single-game MVP evaluation
Team Composition Stable lineups Random 3v3 matchmaking
Data Type Real-world play-by-play stats In-game event logs
The structural comparison between NBA and Dunk City Dynasty (DCD) (TABLE III) reveals three fundamental differences in their MVP evaluation frameworks:
Evaluation Scope: NBA assesses player performance based on cumulative multi-game statistics over an entire season, while DCD focuses exclusively on single-game MVP determinations.
Team Dynamics: NBA features stable team lineups throughout the season, in contrast to DCD’s dynamic random 3v3 matchmaking system where team compositions vary per game.
Data Characteristics: The analysis relies on fundamentally different data sources - traditional play-by-play statistics for NBA versus structured in-game event logs for DCD. This distinction necessitates different methodological approaches for processing and interpreting the respective data types.
These structural differences highlight our framework’s adaptability to both traditional sports analytics and esports environments with distinct evaluation requirements.
IV-B3 Methodological Adaptations
  • For NBA:
    • Uses traditional basketball stats (e.g., DRtg, BPM)
    • Evaluates season-long contributions
  • For DCD:
    • Processes game-specific metrics (e.g., NicePass, Block)
    • Handles unstable team formations via feature-level Shapley values
IV-B4 Why These Differences Are Strengths
  1. 1.
    Generalizability
    • Handles both:
      • Structured, long-term evaluations (NBA)
      • Dynamic, single-game scenarios (DCD)
  2. 2.
    Robustness
    • Validated on:
      • Expert-curated ground truth (NBA)
      • Crowdsourced consensus (DCD)
  3. 3.
    Scalability
    • Same LightGBM architecture works for:
      • Traditional basketball analytics
      • Esports-specific metrics
IV-C Results
IV-C1 NBA Regular Season MVP Results
[…]
TABLE IV: The top-ten-ranked players during the 2023/2024 NBA regular season MVP results.
Rank GT GSv API
Player Player Contribution $\uparrow$ Player Contribution $\uparrow$
1 Nikola Jokić Jaylen Martin 0.995 Luka Dončić 0.335
2 Shai Gilgeous-Alexander Wenyen Gabriel 0.989 Sam Hauser 0.332
3 Luka Dončić Robert Williams 0.980 Victor Wembanyama 0.331
4 Giannis Antetokounmpo Scoot Henderson 0.976 Nikola Jokić 0.329
5 Jalen Brunson Scotty Pippen Jr. 0.974 Al Horford 0.326
6 Jayson Tatum Keyonte George 0.973 Andre Drummond 0.326
7 Anthony Edwards Killian Hayes 0.968 Stephen Curry 0.325
8 Domantas Sabonis Kelly Olynyk 0.966 James Harden 0.325
9 Kevin Durant Darius Bazley 0.965 Chet Holmgren 0.325
10 - Zach Collins 0.965 Domantas Sabonis 0.325
[…]
Player ACRW $\downarrow$ Player ACRA $\downarrow$ Player AC $\uparrow$
1 Nikola Jokić 2.67 Nikola Jokić 5.52 Nikola Jokić 2.78
2 Shai Gilgeous-Alexander 3.02 Kentavious Caldwell-Pope 5.95 Shai Gilgeous-Alexander 2.31
3 Paul George 3.16 Shai Gilgeous-Alexander 6.01 Kentavious Caldwell-Pope 2.11
4 Luka Dončić 3.36 Jayson Tatum 6.49 Jayson Tatum 1.91
5 Bogdan Bogdanović 4.07 Paul George 6.75 Derrick White 1.73
6 Jalen Brunson 4.09 Jusuf Nurkić 6.90 Rudy Gobert 1.66
7 Victor Wembanyama 4.11 Rudy Gobert 7.12 Paul George 1.56
8 Tre Jones 4.19 Derrick White 7.13 Isaiah Hartenstein 1.51
9 Jayson Tatum 4.30 Michael Porter Jr. 7.31 Jusuf Nurkić 1.48
10 Devin Booker 4.34 Isaiah Hartenstein 7.33 Michael Porter Jr. 1.47
[…]
Player ACRW $\downarrow$ Player ACRA $\downarrow$ Player AC $\uparrow$
1 Nikola Jokić 2.38 Nikola Jokić 3.23 Nikola Jokić 7.39
2 Giannis Antetokounmpo 2.97 Giannis Antetokounmpo 3.47 Giannis Antetokounmpo 6.14
3 Shai Gilgeous-Alexander 3.49 Shai Gilgeous-Alexander 4.17 Luka Dončić 5.76
4 Tyrese Haliburton 3.96 Luka Dončić 4.30 Shai Gilgeous-Alexander 5.61
5 Domantas Sabonis 4.01 LeBron James 5.38 Jalen Brunson 4.00
6 Luka Dončić 4.02 Daniel Gafford 5.53 LeBron James 3.91
7 Anthony Davis 4.15 Anthony Davis 5.54 Domantas Sabonis 3.78
8 Daniel Gafford 4.70 Domantas Sabonis 5.67 Anthony Davis 3.72
9 Jalen Brunson 4.93 Isaiah Hartenstein 5.77 Tyrese Haliburton 3.59
10 Victor Wembanyama 5.05 Jarrett Allen 5.81 Kawhi Leonard 3.21
[…]
Player ACRW $\downarrow$ Player ACRA $\downarrow$ Player AC $\uparrow$
1 Victor Wembanyama 2.21 Nikola Jokić 4.85 Nikola Jokić 1.31
2 Nikola Jokić 2.71 Victor Wembanyama 5.23 Victor Wembanyama 0.97
3 Daniel Gafford 2.93 Rudy Gobert 5.73 Rudy Gobert 0.96
4 Rudy Gobert 3.24 Giannis Antetokounmpo 6.36 Shai Gilgeous-Alexander 0.83
5 Nick Richards 3.35 Jusuf Nurkić 6.50 Anthony Davis 0.60
6 Nic Claxton 3.51 Shai Gilgeous-Alexander 6.93 Giannis Antetokounmpo 0.51
7 Giannis Antetokounmpo 3.88 Anthony Davis 6.94 Chet Holmgren 0.48
8 Anthony Davis 3.91 Chet Holmgren 7.04 Bam Adebayo 0.40
9 Bam Adebayo 4.82 Andre Drummond 7.05 Jusuf Nurkić 0.39
10 Zion Williamson 4.85 Daniel Gafford 7.05 Isaiah Hartenstein 0.38
TABLE V: The top-all-ranked players during the 2022-2024 NBA regular season MVP results.
Method 2023/2024 2022/2023 2021/2022
ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$ ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$ ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$
[…]
GSV 228.89 -0.60 0.00 212.84 0.32 0.00 206.75 -0.13 0.00
API 36.64 0.32 0.22 55.00 0.45 0.46 53.67 0.52 0.42
TABLE VI: The top-three-ranked players during the 2022-2024 NBA regular season MVP results.
Method 2023/2024 2022/2023 2021/2022
ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$ ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$ ARD $\downarrow$ SRCC $\uparrow$ R $\uparrow$
[…]
GSV 297.0 -0.50 0.00 169.00 -0.50 0.00 198.00 0.50 0.00
API 22.33 -0.50 0.33 5.33 -0.50 0.33 2.67 0.50 0.33
IV-C2 NBA Finals MVP Results
We computed the NBA’s Finals Most Valuable Players (FMVPs) over the past decade, presenting the detailed results in TABLE VII. We conducted tests using three versions of the win-loss model across three MVP evaluation methods. Subsequently, we documented the corresponding FMVP rankings for each method. A smaller ranking indicates closer proximity to the ground truth, with a ranking of 1 signifying that the MVP evaluated by this method aligns with the FMVP of the year. Our experimental findings demonstrate that the first version of the win-loss model yields the least favorable outcomes, whereas the third version exhibits the highest consistency with the ground truth. Interestingly, this outcome slightly deviates from regular season results, suggesting that there is a difference in how defensive efficiency is considered between regular season and playoff voting.
TABLE VII: NBA finals MVP results.
Method (Rank $\downarrow$) 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015
FMVP (GT) Jaylen Brown Nikola Jokić Stephen Curry Giannis Antetokounmpo LeBron James Kawhi Leonard Kevin Durant Kevin Durant LeBron James Andre Iguodala
[…]
IV-C3 Dunk City Dynasty MVP Results
(a) ARD
(b) SRCC
Figure 3: The dashed horizontal line represents the mean, the red horizontal line represents the median, and the orange triangles are outliers.
TABLE VIII: The Dunk City Dynasty MVP results.
Method ARD $\downarrow$ SRCC $\uparrow$ ACC $\uparrow$
Baseline 0.42 $\pm$ 0.44 0.62 $\pm$ 0.51 0.54
Ours 0.26 $\pm$ 0.33 0.80 $\pm$ 0.24 0.83
MVP results with crowdsourced confidence.
We sourced a total of 500 videos through crowdsourcing and subsequently considered the results with crowdsourced confidence exceeding 0.9 as the golden dataset. We divided the golden dataset into the dataset required for feature optimization and the test set in a 1:1 ratio. Traditional baselines (GSv) are no longer applicable in the Dunk City Dynasty due to the random grouping of players. Dunk City Dynasty currently utilizes an API-based metric-weighted approach for online MVP calculation, which serves as our baseline. Specifically, we utilized the ‘MvpPoint’ from the settlement data of each game as the baseline result. The results of various metrics for Dunk City Dynasty MVP evaluation are presented in TABLE VIII and Fig. 3. Our method demonstrates superior performance compared to the baseline. Furthermore, we explored the influence of crowdsourced confidence on the results. We divided the confidence intervals into 0.0-0.5, 0.5-0.9, 0.9-1.0, and the ratio of the number of games within these three confidence intervals is 1:1:1. As depicted in Fig. 4 and TABLE IX, a decrease in crowdsourced confidence results in a decline in the effectiveness of the MVP evaluation using our method. Furthermore, the performance of the Baseline in the confidence interval of 0.0-0.5 is better than that in 0.5-0.9, indicating poor interpretability.
(a) ARD
(b) SRCC
Figure 4: Mean-variance plot of the evaluation metrics to crowdsourced confidence.
TABLE IX: Results of accuracy changes with crowdsourced confidence interval variations.
Confidence Interval 0.9-1.0 0.5-0.9 0.0-0.5
Baseline (ACC $\uparrow$) 0.54 0.49 0.43
Ours (ACC $\uparrow$) 0.83 0.82 0.76
MVP results with GOSDT tree [27].
However, the golden dataset only accounts for one-fifth of all crowdsourcing results, to fully utilize the voting results from all game videos, we employed the state-of-the-art voting algorithm, the GOSDT tree, to select the MVP for each match. Consequently, every match obtained an MVP ground truth. Ultimately, the obtained ground truth from this algorithm was used to validate the superiority of our approach further. As the voting algorithm selects only one MVP instead of a ranking, we evaluated using two metrics: $ARD$ and $ACC$. The experimental results in TABLE X demonstrate the superiority of our method.
TABLE X: The Dunk City Dynasty MVP results with GOSDT.
Method ARD $\downarrow$ ACC $\uparrow$
Baseline 0.59 $\pm$ 0.65 0.50
Ours 0.35 $\pm$ 0.59 0.72
IV-C4 Ablation Study
We evaluated three method variants as shown in TABLE XI across all experiments:
TABLE XI: Method Variants and Components
Version Components Included
Ours_V1 Base model (no optimizations)
Ours_V2 + Causal inference only
Ours_V3 + Causal + Fuzzification
The comprehensive experimental results are presented across:
  • Quantitative comparisons: TABLE IV-TABLE VI demonstrate performance improvements at each stage
  • Qualitative analysis: Fig. 6-Fig. 7 visualize the feature attribution patterns
Key Findings from Ablation
  1. 1.
    Causal Inference Alone (Ours_V2)
    • 30.9% relative improvement of ARD vs Ours_V1 in TABLE V
    • Improved ranking consistency of SRCC vs Ours_V1 in TABLE VI
    • 19.6% relative improvement of R vs Ours_V1 in TABLE V
  2. 2.
    Fuzzification Added (Ours_V3)
    • Further stabilized defensive metrics’ contributions (Fig. 6)
    • NBA Finals MVP results are more accurate (TABLE VII)
IV-C5 Online Service Deployment Framework for MVP-Shapley
Figure 5: Online Service Deployment Framework.
Fig. 5 illustrates the deployment framework of our algorithm in the online service. Hive and Impala are both widely used technologies in the Hadoop ecosystem, designed for querying and analyzing large datasets stored in Hadoop Distributed File System (HDFS). The MVP evaluation deployment framework consists of three main modules: Data, training, and serving. 1) Data module: Game logs from live basketball games are collected and processed using Hive/Impala. The data, including game ID and player statistics (e.g., point, assist, etc.), is preprocessed to form training data. 2) Training module: The preprocessed data is used in the model training phase to generate the MVP evaluation model. The trained model is stored in a model storage for later use. 3) Serving module: Live game data from the game client is processed and passed to an online inference service. Using the pre-trained model, the service evaluates the MVPs in real time, with results sent back to the client for immediate display. This framework provides a scalable and efficient approach to deploying MVP-Shapley, seamlessly integrating historical data with real-time game information to deliver accurate evaluations during live matches.
For the performance comparison with the original method, we conducted an online A/B test for over one month to compare player report rates ($RR=\frac{N_{\text{reports}}}{N_{\text{players}}}\times 100\%$) and churn rates ($CR=\frac{N_{\text{churned}}}{N_{\text{start\_players}}}\times 100\%$). Where $N_{\text{reports}}$ = Number of reports received, $N_{\text{players}}$ = Total number of active players during the same period, $N_{\text{churned}}$ = Number of players who stopped playing the game, $N_{\text{start\_players}}$ = Total number of active players at the beginning of the period. We employed a rigorous hash-based random bucketing method for our A/B test, with 10,000 players assigned to each bucket. Our method has reduced the report rate by approximately $9.64\text{\textperthousand}\pm 1.2\text{\textperthousand}$ and the churn rate by approximately $8.11\text{\textperthousand}\pm 0.9\text{\textperthousand}$ compared to the existing method. Experimental results indicate the superiority of our method.
IV-C6 Computational Efficiency of Shapley Values
We leverage TreeSHAP [23] to compute feature-level Shapley values directly from the LightGBM model, which reduces complexity from $\mathcal{O}(TL2^{M})$ to $\mathcal{O}(TLD^{2})$ (where $T$ = number of trees, $L$ = max leaves, $D$ = depth, $M$ = features) and enables linear-time computation relative to tree size.
Our experiments ran on:
  • CPU: 4 cores of AMD EPYC 7543 (32-core)
  • GPU: NVIDIA Tesla A30 (24GB VRAM)
Achieving:
  • Sub-second response for single-game MVP evaluation
  • <<5 minutes for full season NBA analysis
IV-D Feature Visualization Analysis
Fig. 6 and Fig. 7 respectively demonstrate the change in feature importance before and after optimization for the NBA dataset and the Dunk City Dynasty dataset. Fig. 6 shows the feature importance without any processing, where the ‘+/-’ feature dominates the results, making it difficult for other features to contribute to the interpretation—an unreasonable scenario. After removing such features and applying fuzzification, Fig. 6 reveals a more balanced influence of different features on the results, enabling each feature to contribute meaningfully to the interpretation. Fig. 8 illustrates the impact of different features on the outcome of actual games for Dunk City Dynasty. These results can help players analyze their performance and serve as a personalized coach to improve specific skills.
Figure 6: NBA dataset. (a) The dot chart visualization of SHAP (SHapley Additive exPlanations) values of all features; (b) The dot chart visualization of SHAP values of the features after fuzzification.
Figure 7: Dunk City Dynasty dataset. (a) The dot chart visualization of SHAP values of all features; (b) The dot chart visualization of SHAP values of optimized features.
Figure 8: Dunk City Dynasty dataset. (a) A case analysis of a home team’s victory; (b)A case analysis of a home team’s loss.
V Related Work
Related work can be categorized into three types: Metric Weighting, Machine Learning Techniques, and Cooperative Game Theory among Players.
1) Metric weighting encompass a variety of single-metric evaluations. For example, PM stands for "Plus-Minus", which measures the difference between the points scored and the points lost by a team when the player is on the court. APM [3, 4] stands for "Adjusted Plus-Minus", a statistic used to evaluate the impact of a player on the team’s score, and RPM [7] is a regularized APM. BPM [5, 6] stands for "Box Plus/Minus", which is used to evaluate a player’s overall efficiency, considering the player’s impact in scoring, rebounding, etc. WS stands for "Win Shares", a metric used to measure a player’s contribution to his team’s wins. WS48 [8] stands for "Win Shares per 48 minutes", which considers a player’s playing time, making comparisons between different players fairer. WARP finds the player’s contribution in terms of how many additional wins he/she brings to the team. VORP [9] is a statistic that combines the strengths of BPM and WARP to estimate the points a player contributes per 100 team possessions above that of a replacement-level player. These metrics can be used to evaluate the MVP individually or in a weighted manner. The weighting of different metrics was introduced early on [10, 11].
2) Machine learning techniques. Fearnhead et al. [12] introduce a new model for evaluating NBA player abilities by comparing team performance with and without the player, controlling for teammates’ abilities. The method uses multi-season data to estimate offensive and defensive capabilities, providing an overall player rating. Page et al. [13] model basketball player performance using Gaussian process regression, estimating player performance curves as a function of game percentile. Metulini et al. [14] emphasize evaluating player performances in team sports, estimating scoring probabilities, and developing player-specific shooting performance indices using Classification and Regression Trees (CART) and game data. Sandri et al. [15] focus on modeling shooting performance variability and teammate interactions using a Markov switching model, highlighting positive and negative interactions between teammates through network graphs. Terner et al. [16] explore various tools for assessing players and discuss the future of basketball analytics, emphasizing the need for causal inference in sports.
3) Cooperative Game Theory among Players. Cooperative game theory, inspired by previous work [17, 18], led Metulini et al. [19] to assess players’ importance in basketball using the generalized Shapley value. They employed various lineup win-rate prediction models based on historical team roster data to compute the MVP for each team and determine the best lineup.
VI Conclusion and Future Work
In this work, we propose MVP-Shapley for evaluating the MVP by leveraging feature-level Shapley values. Our method is both scalable and interpretable. MVP-Shapley breaks down the contribution of players to the game victory by attributing it to each player’s features, then calculates the Shapley value for each feature. We validated our approach on the NBA and Dunk City Dynasty datasets and successfully deployed it for online industrial use. Future work will focus on enhancing the method’s robustness and stability for broader practical applications. Additionally, we aim to explore finer-grained tracking data, such as player state-action time series, to improve interpretability and uncover deeper insights into player performance.